OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

-6 , 4.5, 6, 14 are options for the answer

OpenStudy (anonymous):

substitute x = -5 (╯° Д° )╯︵ ┻━┻

OpenStudy (anonymous):

wat???

OpenStudy (tyteen4a03):

Substitute -5 into the expression, you have −2(-5) + 4. Now do the math and you should get the answer.

OpenStudy (anonymous):

g(-5) = −2(-5) + 4 can you solve this^?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

why not :( ? I can't give answers. let's go solve step by step: −2(-5) + 4 -2 times -5 = ?

OpenStudy (anonymous):

okay plzz tell me n ik yall cant give out the answer

OpenStudy (anonymous):

\[-2 \times -5 = ?\]do this first, you know multiplication, no?

OpenStudy (anonymous):

umm 10

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

then 10 + 4 = ?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

so (-2)(-5) + 4 = 14 there u go lol

OpenStudy (anonymous):

yhxs so thats the answer lol

OpenStudy (anonymous):

right?